# Tianjun Gu1 & Linfeng Li1 & Xuhong Wang2 & Chenghua Gong1 & Jingyu Gong1 & Zhizhong Zhang1 & Yuan Xie1,3 & Lizhuang Ma1 & Xin Tan1,2

### v3

@@ Abstract
Adaptive navigation in unfamiliar environments is crucial for household service robots but remains challenging due to the need for both low-level path planning and high-level scene understanding. While recent vision-language model (VLM) based zero-shot approaches reduce dependence on prior maps and scene-specific training data, they face significant limitations: spatiotemporal discontinuity from discrete observations, unstructured memory representations, and insufficient task understanding leading to navigation failures. We propose DORAEMON (Decentralized Ontology-aware Reliable Agent with Enhanced Memory Oriented Navigation), a novel cognitive-inspired framework consisting of Ventral and Dorsal Streams that mimics human navigation capabilities. The Dorsal Stream implements the Hierarchical Semantic-Spatial Fusion and Topology Map to handle spatiotemporal discontinuities, while the Ventral Stream combines RAG-VLM and Policy-VLM to improve decision-making. Our approach also develops Nav-Ensurance to ensure navigation safety and efficiency. We evaluate DORAEMON on the HM3D, MP3D, and GOAT datasets, where it achieves state-of-the-art performance on both success rate (SR) and success weighted by path length (SPL) metrics, significantly outperforming existing methods. We also introduce a new evaluation metric (AORI) to assess navigation intelligence better. Comprehensive experiments demonstrate DORAEMON’s effectiveness in zero-shot autonomous navigation without requiring prior map building or pre-training.
@@
footnotetext: Homepage: https://grady10086.github.io/DORAEMON/
@@ 1 Introduction
Adaptive navigation in complex and unseen environments [2] is a key capability for household service robots. This task requires robots to move from a random starting point to the location of a target object without any prior knowledge of the environment. For humans, navigation appears almost trivial thanks to spatial cognitive abilities honed through evolution. For robots, however, navigation remains a highly challenging problem: it demands not only low-level path planning to avoid obstacles and reach the destination, but also high-level scene understanding to interpret and make sense of the surrounding environment.
@@
Most existing navigation methods rely on the construction of prior maps[4] or require extensive scene-specific data for task-oriented pre-training[36]. However, in entirely unfamiliar environments, building maps in advance can be time-consuming and labor-intensive, and the scarcity of scene data further limits the practicality of these approaches. Recently, some works [14, 34] have begun to explore zero-training and zero-shot navigation strategies based on vision-language models (VLMs). Relying on textual descriptions of the current task, image inputs, and previously observed historical information, these approaches achieve navigation without dependence on environment- or task-specific data, gradually shedding the reliance on scene priors.
@@
Although VLM-based zero-shot navigation methods offer a novel perspective for adapting to unfamiliar environments, they still face numerous challenges in practical applications. On the one hand, VLMs typically take task descriptions and observation histories as input [34][14]. However, due to the discrete nature of input image descriptions at each time step, this spatiotemporal discontinuity often makes it difficult for VLMs to understand the relationships between targets and obstacles in complex environments. On the other hand, while many existing navigation systems incorporate some form of memory functionality, most VLM methods [30, 25, 27] adopt a single-step decision paradigm, treating historical information merely as a reference log. Even though End-to-End methods like VLMnav[14] utilize historical information, they typically store this information in a flat, unstructured manner, which fundamentally limits their ability to perform long-range navigation. Additionally, VLMs sometimes insufficient understanding of task semantics often leads to poor decision-making (e.g., going to look for a TV but finding a computer instead), and the lack of reliable check mechanisms for navigation states frequently results in unreliable behaviors such as spinning in place during navigation tasks. Fig 1 conceptually illustrates limitations of traditional VLN methods and contrasts them with the cognitive-inspired approach of our DORAEMON.
Figure 1: (a) Illustrates limitation of typical VLM navigation (red arrow) in an unfamiliar environment. (b) DORAEMON’s cognitive inspiration: a Decentralized Ontology-aware approach with a Dorsal Stream for “Where” and a Ventral Stream for “What”. (c) DORAEMON constructs the Topology Map and uses Hierarchical Semantic-Spatial Fusion(green arrow).
Inspired by cognitive science ‘Decentralized Ontology’ principles [3] suggesting human knowledge is often represented and accessed in a distributed and context-dependent manner rather than via a single monolithic structure, we propose the Decentralized Ontology-aware Reliable Agent with Enhanced Memory Oriented Navigation (DORAEMON), which consists mainly of Ventral Stream and Dorsal Stream. The Ventral Stream processes object identity (“what”) information, while the Dorsal Stream handles spatial (“where”) processing in the human brain.
@@
Our decentralized architecture distributes cognitive functions across complementary streams: The Dorsal Stream addresses spatio-temporal discontinuities through a Topology Map and a Hierarchical Semantic-Spatial Fusion, allowing our agent to reason accurately about target-environment relationships. Additionally, the Ventral Stream improves task understanding by utilizing a Retrieval-augmented Generation model (RAG-VLM) and Policy-VLM for navigation. Additionally, DORAEMON features a Nav-Ensurance system that enables agents to autonomously detect and respond to abnormal conditions, such as becoming stuck or blocked during navigation. To evaluate navigation performance more comprehensively, we propose a new metric called the Adaptive Online Route Index (AORI).
@@
  • We propose DORAEMON, a novel adaptive navigation framework inspired by cognitive principles of decentralized knowledge, consisting of ventral and Dorsal Streams, enabling End-to-End and zero-shot navigation in completely unfamiliar environments without pre-training.
  • We propose the Dorsal Stream, which involves designing a Topology Map and a Hierarchical Semantic-Spatial Fusion Network to effectively manage spatio-temporal discontinuities. Additionally, we introduce the Ventral Stream, incorporating a synergistic reasoning component that combines RAG-VLM for understanding ontological tasks and Policy-VLM for enhanced task comprehension and policy planning.
@@ 2 Related Work
2.1 Zero-shot Object Goal Navigation
Object navigation methods are broadly supervised or zero-shot. Supervised approaches train visual encoders with reinforcement/imitation learning [23, 17, 31, 11] or build semantic maps from training data [45, 24, 49], struggling with novel scenarios due to data dependency. Zero-shot methods address this using open-vocabulary understanding, increasingly leveraging foundation models like Vision-Language Models (VLMs) and Large Language Models (LLMs). LLMs provide commonsense reasoning via object-room correlation [41, 51, 39], semantic mapping [43], and chain-of-thought planning [6, 41, 33, 5], while VLMs align visual observations with textual goals. These foundation model-guided techniques include image-based methods mapping targets to visual embeddings [38, 22, 13, 1] and map-based approaches using frontier [50, 45, 51, 10, 19, 43, 33] or waypoint-based maps [39] with LLM/VLM reasoning. VLM-based strategies either use VLMs for recognition with traditional planning and extra perception models [27, 46, 21, 44, 37], or, like PIVOT [25] and VLMnav [14], directly produce actions End-to-End via visual prompting. Despite progress, many zero-shot methods, especially those processing observations independently, face challenges integrating temporal information and handling complex spatial reasoning in unfamiliar environments.
@@ 2.2 Memory Mechanisms in Navigation
Memory representations in navigation systems have evolved through various architectures, including episodic buffers that maintain observation sequences [14, 34, 16], spatial representations prioritizing geometric information [50, 46], graph-based semantic structures capturing object relationships [41], and predictive world models attempting to forecast environmental states [7, 26]. These systems typically process semantic and spatial information separately, with limited integration between perception and reasoning modules. Most approaches focus on either building representations or enhancing reasoning mechanisms independently. Differently, DORAEMON integrates these aspects through a hierarchical semantic-spatial fusion network with bidirectional information flow between ventral and dorsal processing streams.
@@ 2.3 Cognitive Neuroscience Inspiration in Navigation
Object navigation systems are influenced by cognitive neuroscience, especially Decentralized Ontology[3], which suggests that human knowledge is organized through interconnected cognitive systems that enable context-dependent reasoning. Recent models like CogNav[7] and BrainNav[20] incorporate cognitive elements, but they do not fully embody Decentralized Ontology. CogNav utilizes a finite state machine for cognitive states, but may have limitations in knowledge integration. BrainNav mimics biological functions but doesn’t deeply engage in decentralized information processing.
In contrast, DORAEMON aims for a more comprehensive Decentralized Ontology-aware processing. It emphasizes the integration and bidirectional exchange of information between Dorsal Stream and Ventral Stream, allowing for the construction of semantic relationships that enhance spatial understanding and support flexible, context-aware navigation.
@@ 3 Methods
We address the ObjectNav task [2], where an agent, starting from an initial pose, must locate and navigate to a target object within a previously unseen indoor environment. At step $t$, the agent receives observation $I_{t}$, current pose $P_{t}$ and a task specification $T$, which can be either a simple object category (e.g., “sofa”) or an instruction (e.g., “find the red chair” or“the plant on the desk”) for tasks like GOAT [18]. Based on these inputs, the agent must decide on an action $a_{t}$. While many prior works utilize a discrete action space, our End-to-End framework employs a continuous action representation in polar coordinates $(r_{t},\theta_{t})$, where $r_{t}$ specifies the forward distance to move, and $\theta_{t}$ denotes the change in orientation. Crucially, the action space also includes a stop action. The task is considered successful if the agent executes the stop action after meeting successive stop triggers in steps $t$ and $t+1$. The trigger occurs when 1) the agent is within a predefined distance threshold $d_{success}$ of the target object; 2) the target object is visually confirmed within the agent’s current observation $I_{t}$.
@@
Our DORAEMON framework achieves End-to-End and zero-shot navigation through the ontology of two decentralized cognitive-inspired streams, as depicted in Figure 2. Given an input with a panoramic image $I_{t}$ and a pose $P_{t}$ at step $t$, they are processed by the Action Proposer Module (Appendix A) and the Dorsal Stream Module (Section 3.1), respectively. In the Action Proposer Module, a candidate image $I^{t}_{\text{anno}}$ is generated with a set of action candidates $A_{\text{final}}^{t}$ through parameterized geometric reasoning. Concurrently, the Dorsal Stream extracts semantic and spatial information from $I_{t}$ using Hierarchical Semantic-Spatial Fusion and stores it within the Topology Map as node ${v}_{t}$. The relevant node ${v}_{\text{rele}}$ can be accessed by op-down retrieval. After that, ${v}_{\text{rele}}$ and $I^{t}_{\text{anno}}$ are input to the Policy-VLM to select the best action based on the given information(Section 3.2.2). At the same time, the Policy-VLM receives a database containing information on several key dimensions relevant to the navigation task $T$, which is generated by the RAG-VLM (Section3.2.1) in the Ventral Stream Module (Section 3.2). The Policy-VLM integrates the information through a chain of thought (Appendix H), identifies abnormal conditions (Section 3.3), and outputs the final choice action $a_{t}$. The agent performs this action $a_{t}$ in the environment, navigates, and makes the next decision at step $t+1$.
@@ 3.1.1 Topological Map
The topological map, defined as $\mathcal{G}=(\mathcal{V},\mathcal{E})$, represents the environment and incorporates historical observations. This map is built incrementally as the agent explores. Each node $v_{t}\in\mathcal{V}$ in the topological map formally integrates multimodal observations as:
$v_{t}=\left(p_{t},q_{t},I_{t},L_{t},o_{t},\mathbf{s}_{t}\right),$ (1)
@@
where $p_{t}$, $q_{t}$, $I_{t}$, $L_{t}$, $o_{t}$, $s_{t}$ correspond to the agent’s position, orientation represented as unit quaternion from agent pose $P_{t}$, visual observation, language description of $I_{t}$, target likelihood estimation, and optional semantic embedding (e.g., CLIP features).The topological map $\mathcal{G}$ is built progressively through spatio-temporal criteria. A new node $v_{\text{new}}$ is added to the node set $\mathcal{V}$ if either of the following conditions is met: 1) Temporal Criterion: A fixed number of exploration steps $S_{\text{update}}$ have elapsed since the last node addition: $t_{\text{curr}}-t_{\text{prev}}\geq S_{\text{update}}$, ensuring periodic state updates; 2) Spatial Criterion: The agent’s displacement from the last node exceeds a spatial sampling threshold: $\|{p}_{\text{curr}}-{p}_{\text{prev}}\|_{2}>\delta_{\text{sample}},$. where $\delta_{\text{sample}}$ controls exploration granularity. When $v_{\text{new}}$ is added to $\mathcal{V}$, it is immediately connected to the nearest node $v_{\text{prev}}$.
@@ 3.1.2 Hierarchical Semantic-Spatial Fusion
Hierarchical Memory Retrieval. To efficiently find relevant information within the constructed hierarchy (e.g., observations related to sofa), the system employs a top-down search, conceptually outlined in the AppendixG. This search is guided by a scoring function $S(n_{i})$ evaluated at nodes $h_{i}$ during traversal the constructed hierarchy:
$S(h_{i})=\alpha_{\text{sem}}S_{\text{semantic}}(h_{i},T)+\alpha_{\text{spa}}S_{\text{spatial}}(h_{i})+\alpha_{\text{key}}S_{\text{keyword}}(h_{i},T)+\alpha_{\text{time}}S_{\text{time}}(h_{i}),$ (3)
where $S_{\text{semantic}}$ computes embedding similarity between node $n_{i}$ and task $T$, $S_{\text{spatial}}$ measures proximity to current position, $S_{\text{keyword}}$ evaluates keyword overlap, and $S_{\text{time}}$ prioritizes recent observations. The weights $\alpha$ balance these components based on their relative importance. To manage computational cost, the retrieval process incorporates beam search, expanding only the top-scoring nodes at each level.

### v4

@@ Abstract
Adaptive navigation in unfamiliar environments is crucial for household service robots but remains challenging due to the need for both low-level path planning and high-level scene understanding. While recent vision-language model (VLM) based zero-shot approaches reduce dependence on prior maps and scene-specific training data, they face significant limitations: spatiotemporal discontinuity from discrete observations, unstructured memory representations, and insufficient task understanding leading to navigation failures. We propose DORAEMON (Decentralized Ontology-aware Reliable Agent with Enhanced Memory Oriented Navigation), a novel cognitive-inspired, zero-shot, end-to-end framework consisting of Ventral and Dorsal Streams that mimics human navigation capabilities. The Dorsal Stream implements the Hierarchical Semantic-Spatial Fusion and Topology Map to handle spatiotemporal discontinuities, while the Ventral Stream combines CoDe-VLM and Exec-VLM to improve decision-making. Our approach also develops Nav-Ensurance to ensure navigation safety and efficiency. We evaluate DORAEMON on the HM3Dv1, HM3Dv2, MP3D, where it achieves state-of-the-art performance on both SR and SPL metrics, significantly outperforming existing methods. We also introduce a new evaluation metric (AORI) to assess navigation intelligence better. Comprehensive experiments demonstrate DORAEMON’s effectiveness in zero-shot and end-to-end navigation without requiring prior map building or pre-training. Our code is available at https://grady10086.github.io/DORAEMON/.
@@ 1 Introduction
Figure 1: Performance comparison of end-to-end methods in object navigation(ob-nav) and multi-modal lifelong navigation(mml).
@@
Adaptive navigation in complex and unseen environments [2] is a key capability for household service robots. This task requires robots to move from a random starting point to the location of a target object without any prior knowledge of the environment. For humans, navigation appears almost trivial, however,navigation remains a highly challenging problem for robots: it demands not only low-level path planning to avoid obstacles and reach the destination, but also high-level scene understanding to interpret and make sense of the surrounding environment.
@@
Most existing navigation methods rely on the construction of prior maps[4] or require extensive scene-specific data for task-oriented pre-training[35]. Recently, some works[39, 48, 37] have begun to explore zero-training and zero-shot navigation strategies, relying on textual descriptions of the current task, image inputs, and previously observed historical information, these approaches achieve navigation without dependence on environment or task-specific data, gradually shedding the reliance on scene priors.
Although zero-shot and zero-training navigation methods offer a novel perspective, they still face numerous challenges in practical applications. On the one hand, most current navigation methods are non-end-to-end, where the agent’s spatial actions is mapped to a discrete set. These discrete actions result in paths that are neither smooth nor efficient. To align with a target, the agent may require multiple small-angle rotations. On the other hand, the primary bottleneck for current Vision-Language Models (VLMs) in long-range navigation is their inadequate memory mechanisms. Their reliance on discrete observational inputs prevents a cohesive understanding of spatiotemporal continuity. More critically, the prevalent approaches[29, 24] of storing history as an unstructured log within a single-step decision paradigm fundamentally compromises their ability to perform effective long-term path planning.
@@
Even though end-to-end methods like VLMnav[24, 14] utilize historical information, they typically store this information in a flat, unstructured manner, which fundamentally limits their ability to perform long-range navigation. Additionally, VLMs sometimes insufficient understanding of task semantics often leads to poor decision-making, and the lack of reliable check mechanisms for navigation states frequently results in unreliable behaviors such as spinning in place during navigation tasks.
@@
Inspired by cognitive science "Decentralized Ontology" principles [3], we propose the Decentralized Ontology-aware Reliable Agent with Enhanced Memory Oriented Navigation (DORAEMON), which consists mainly of a Ventral Stream and a Dorsal Stream. The core theoretical premise is that knowledge is inherently distributed and context-dependent, composed of multiple local perspectives, rather than the single, monolithic world model. The Ventral Stream processes object identity (“what”) information, while the Dorsal Stream handles spatial (“where”) processing in the human brain. The Dorsal Stream addresses spatio-temporal discontinuities through a Topology Map and a Hierarchical Semantic-Spatial Fusion, allowing our agent to reason accurately about target-environment relationships. Additionally, the Ventral Stream improves task understanding by utilizing a CoDe-VLM (Compositional Decomposition VLM) and Exec-VLM (Execution VLM) for navigation. Additionally, DORAEMON features a Nav-Ensurance system that enables agent to autonomously detect and respond to abnormal conditions, such as becoming stuck or blocked during navigation. To evaluate navigation performance more comprehensively, we propose a new metric called the Adaptive Online Route Index (AORI). Fig 2 conceptually illustrates limitations of traditional VLN methods and contrasts them with DORAEMON.
@@
  • We propose DORAEMON, a novel adaptive navigation framework inspired by cognitive principles of decentralized knowledge, consisting of ventral and Dorsal Streams, enabling end-to-end and zero-shot navigation in completely unfamiliar environments without pre-training, offering plug-and-play compatibility with any VLMs.
  • We propose the Dorsal Stream, which involves designing a Topology Map and a Hierarchical Semantic-Spatial Fusion Network to effectively manage spatio-temporal discontinuities. Additionally, we introduce the Ventral Stream, incorporating a synergistic reasoning component that combines CoDe-VLM for understanding ontological tasks and Exec-VLM for enhanced task comprehension and planning.
@@
Figure 2: (a) Illustrates limitation of typical VLM navigation (red arrow). (b) DORAEMON’s cognitive inspiration from human navigation. (c) Our DORAEMON method.
@@ 2 Related Work
2.1 Zero-shot Navigation
Navigation methods are broadly supervised or zero-shot. Supervised approaches train visual encoders with reinforcement/imitation learning [16, 22, 30, 11] or build semantic maps from training data [43, 23, 47], struggling with novel scenarios due to data dependency. Zero-shot methods address this using open-vocabulary understanding, increasingly leveraging foundation models like LLMs and VLMs. LLMs provide commonsense reasoning via object-room correlation [39, 49, 37], semantic mapping [42], and chain-of-thought planning [6, 39, 32], while VLMs align visual observations with textual goals. These foundation model-guided techniques include image-based methods mapping targets to visual embeddings [36, 13, 1] and map-based approaches using frontier [48, 43, 10, 18, 42, 32] or waypoint-based maps [37] with LLM/VLM reasoning. VLM-based strategies either use VLMs for recognition with traditional planning and extra perception models [26, 44], or, like PIVOT [24] and VLMnav [14], directly produce actions end-to-end via visual prompting. Despite progress, many zero-shot methods, especially those processing observations independently, face challenges integrating temporal information and handling complex spatial reasoning in unfamiliar environments.
@@ 2.2 Memory Mechanisms in Navigation
Memory representations in navigation systems have evolved through various architectures, including episodic buffers that maintain observation sequences [14, 33, 15], spatial representations prioritizing geometric information [48, 44], graph-based semantic structures capturing object relationships [40, 39], predictive world models attempting to forecast environmental states [7, 25] and the memory capacity acquired through training[50]. These systems typically process semantic and spatial information separately, with limited integration between perception and reasoning modules. Most approaches focus on either building representations or enhancing reasoning mechanisms independently. Differently, DORAEMON integrates these aspects through a hierarchical semantic-spatial fusion network with bidirectional information flow between ventral and dorsal processing streams.
@@ 2.3 Cognitive Neuroscience Inspiration in Navigation
Navigation systems are influenced by cognitive neuroscience, recent models like CogNav[7] and BrainNav[19] incorporate cognitive elements, but they do not fully embody Decentralized Ontology. CogNav utilizes a finite state machine for cognitive states, but may have limitations in knowledge integration. BrainNav mimics biological functions but doesn’t deeply engage in decentralized information processing. In contrast, DORAEMON is inspired by Decentralized Ontology[3], which suggests that human knowledge is organized through interconnected cognitive systems that enable context-dependent reasoning. It emphasizes the integration and bidirectional exchange of information between Dorsal Stream and Ventral Stream, allowing for the construction of semantic relationships that enhance spatial understanding and support flexible, context-aware navigation.
@@ 3 Methods
We address the Navigation task [2], where an agent, starting from an initial pose, must locate and navigate to a target within a previously unseen indoor environment. At step $t$, the agent receives observation $I_{t}$, current pose $P_{t}$ and a task specification $T$, which can be either a simple object category (e.g., “sofa”) or an instruction (e.g., “find the red chair” or“the plant on the desk”) for tasks like GOAT [17]. Based on these inputs, the agent must decide on an action $a_{t}$. While many prior works utilize a discrete action space, our end-to-end framework employs a continuous action representation in polar coordinates $(r_{t},\theta_{t})$, where $r_{t}$ specifies the forward distance to move, and $\theta_{t}$ denotes the change in orientation. Crucially, the action space also includes a stop action. The task is considered successful if the agent executes the stop action after meeting successive stop triggers in steps $t$ and $t+1$. The trigger occurs when 1) the agent is within a predefined distance threshold $d_{success}$ of the target object; 2) the target object is visually confirmed within the agent’s current observation $I_{t}$.
@@
Our DORAEMON framework achieves end-to-end and zero-shot navigation through two decentralized cognitive-inspired streams, as depicted in Figure 3. Given an input with a panoramic image $I_{t}$ and a pose $P_{t}$ at step $t$, they are processed by Action Proposer (Appendix A) and Dorsal Stream(Section 3.1), respectively. In the Action Proposer, a candidate image $I^{t}_{\text{anno}}$ is generated with a set of action candidates $A_{\text{final}}^{t}$. Concurrently, the Dorsal Stream extracts semantic and spatial information from $I_{t}$ using Hierarchical Semantic-Spatial Fusion and stores it within the Topology Map as node ${v}_{t}$. The relevant node ${v}_{\text{rele}}$ can be accessed by up-down retrieval. After that, ${v}_{\text{rele}}$ and $I^{t}_{\text{anno}}$ are input to the Exec-VLM to select the best action based on the given information(Section 3.2.2). At the same time, the Exec-VLM receives a task-specific knowledge graph (KG) relevant to the task $T$, which is generated by the CoDe-VLM (Section3.2.1) in the Ventral Stream (Section 3.2). The Exec-VLM integrates the information through chain of thought (Appendix H), identifies abnormal conditions (Section 3.3), and outputs the final action $a_{t}$. The agent performs this action $a_{t}$ in the environment, navigates, and makes the next decision at step $t+1$.
@@ 3.1.1 Topological Map
The topological map $\mathcal{G}=(\mathcal{V},\mathcal{E})$ is constructed incrementally. Each node $v_{t}\in\mathcal{V}$ encapsulates the agent’s state at timestep $t$, defined as a tuple $(p_{t},q_{t},I_{t},L_{t},o_{t},s_{t})$. $p_{t}$ and $q_{t}$ denote the agent’s position and orientation, which constitute the pose $P_{t}$. $I_{t}$ is the visual observation, $L_{t}$ is its corresponding language description, $o_{t}$ is the target likelihood estimation, and $s_{t}$ represents a semantic embedding of the observation (e.g., CLIP features). A new node $v_{\text{new}}$ is added to $\mathcal{V}$ based on spatio and temporal criteria: a new node is created if either the time elapsed since the last node addition $t_{\text{curr}}-t_{\text{prev}}$ exceeds a temporal threshold $S_{\text{update}}$, or if the agent’s Euclidean distance from the previous node $\|{p}_{\text{curr}}-{p}_{\text{prev}}\|_{2}$ surpasses a spatial threshold $\delta_{\text{sample}}$. Upon its creation, $v_{\text{new}}$ is connected to its predecessor node $v_{\text{prev}}$ via a new edge.
@@ 3.1.2 Hierarchical Semantic-Spatial Fusion
Hierarchical Memory Retrieval. To efficiently find relevant information within the constructed hierarchy (e.g., observations related to sofa), the system employs a top-down search. This search is guided by a scoring function $S(h_{i})$ evaluated at nodes $h_{i}$ during traversal the constructed hierarchy:
$S(h_{i})=\alpha_{\text{sem}}S_{\text{sem}}(h_{i},T)+\alpha_{\text{spa}}S_{\text{spa}}(h_{i})+\alpha_{\text{key}}S_{\text{key}}(h_{i},T)+\alpha_{\text{time}}S_{\text{time}}(h_{i}),$ (2)
where $S_{\text{sem}}$ computes embedding similarity between node $h_{i}$ and task $T$, $S_{\text{spa}}$ measures proximity to the current position using an exponential decay function, $S_{\text{key}}$ evaluates keyword overlap, and $S_{\text{time}}$ prioritizes recent observations. For instance, semantic similarity is calculated via normalized cosine similarity, while spatial and temporal scores both rely on exponential decay models to reflect their diminishing influence over distance and time. (see more details in Appendix G).